Find the cost per ounce of a sunscreen made from 24 ounces of a lotion that costs $1.90 per ounce and 50 ounces of a lotion that costs $3.25 per ounce.
1
Expert's answer
2012-07-19T09:26:32-0400
Let's make an estimation. The final price of a sunscreen is

P = 24*$1.90 + 50*$3.25 = $45.6 + $162.5& = $208.1.

The weight of a sunscreen is

W = 24 + 50 = 74 ounces.

So, the cost per ounce of a sunscreen is

U = P/W = $208.1/74 = $2.812(162) per ounce.